OpenStudy (anonymous):

Solve for x: 1 over 2 plus 4 over 2 x equals quantity x plus 4 over 10

OpenStudy (anonymous):

\[1/2 + 4/2x = x + 4/10\]

OpenStudy (mertsj):

The common denominator is 10x so multiply every term by 10x. That will cancel all denominators.

OpenStudy (anonymous):

\[\frac12 + 4/2x = x + 2/10\] Find a common denominator, 10x \[5x/10x + 20/10x = 10x^{2}/10x + 2x/10x\] Since it's common denominator, just solve what's in the numerator. \[5x + 20 = 10x^2 + 2x\] \[0 = 10x^2 - 3x - 20\] Just solve the quadratic now to get x.

OpenStudy (anonymous):

is it 5?

OpenStudy (anonymous):

No.

OpenStudy (anonymous):

-4?

OpenStudy (anonymous):

Plug it into the quadratic formula. \[-b ± \sqrt{b^{2} - 4ac}\over2a\]

OpenStudy (mertsj):

Calcmathlete gave you wrong information. \[\frac{1}{2}\times10x+\frac{4}{2x}\times10x=\frac{x+4}{10}\times10x\]

OpenStudy (mertsj):

\[5x+20=x^2+4x\]

OpenStudy (mertsj):

\[x^2-x-20=0\]

OpenStudy (mertsj):

\[(x-5)(x+4)=0\]

OpenStudy (mertsj):

\[x=5, -4\]

OpenStudy (anonymous):

Oh. I misread the question. She put 1/2+4/2x=x+4/10, so I thought that somethings were moved around.

OpenStudy (mertsj):

In her first post it says quantity x+4 over 10